Artemis shoots an arrow at a target that is a horizontal distance d=65 m away: the bull's eye of the target is at same height of the arrow. At what angle in degrees above the horizontal must the arrow be released to hit the bull's eye if the arrows initial speed in 39 m/s?
Answer #1

using the equation of linear equation we have

Y = Uyt - 0.5gt^2

= u sin \theta * t- 0.5gt^2

t = 2u sin \theta / g

X = Ux*t

X = u cos \theta * t

X = u*cos \theta * 2u sin \theta / g

X = u^2 sin 2\theta / g

65 = 39^2* sin 2\theta / 9.8

\theta = 12.38o

so the angle of projection = 12.38o